# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

### v1

@@ 4.7 Comparison between the global and IGM escape fractions
Figure 12: Comparison between stacked ${\rm Ly}\alpha$ line profiles. The observed stacked ${\rm Ly}\alpha$ line profile is shown in grey, while zELDA’s prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The stacked ${\rm Ly}\alpha$ line profile is computed in 6 redshift bins. Each panel show the redshift bin indicated in their panel.
In this work, we have inferred the IGM escape fraction,$f_{\rm esc}^{4\AA}$, from the shape of the ${\rm Ly}\alpha$ line profiles. The global ${\rm Ly}\alpha$ escape fraction should always be lower than the ISM, CG,M and IGM escape fractions, as $f_{\rm esc}^{\rm Lya}$ is the product of these. Our direct measurements of $\langle f_{\rm esc}^{4\AA}\rangle$ (blue and green squares for IGM+z and IGM-z) are always above the measured $f_{\rm esc}^{\rm Lya}$ (grey dots), which is reassuring.
Interestingly, we find that the IGM escape fraction decreases with redshift. This is expected from a theoretical point of view given the following two facts. First, it is found by Hayes et al. (2011) and this work that the stacked galactic ${\rm Ly}\alpha$ line profile does not evolve significantly from redshift 0 to 6. Second, the IGM becomes denser at higher redshifts and absorbs more ${\rm Ly}\alpha$ photons (Faucher-Giguère et al., 2008; Laursen et al., 2011b; Gurung-López et al., 2020; Byrohl & Gronke, 2020). Moreover, we have included zELDA’s prediction on observed spectrum at $z>5$ as colored empty squares. $f_{\rm esc}^{4\AA}$ values for $z>5$ sources should be taken cautiously as the IGM transmission curves only reach $z=5.0$. We find that zELDA predicts that $\langle f_{\rm esc}^{4\AA}\rangle$ remains constant from $z=5$ to 6. This can happen since the blue peak of the ${\rm Ly}\alpha$ line is always absorbed, but the red peak is relatively little attenuated. The flattening in $\langle f_{\rm esc}^{4\AA}\rangle$ is also found for $z>4.0$ in the training set of IGM+z (see ZP24), but in the observed data IGM+z only predicts a stable $\langle f_{\rm esc}^{4\AA}\rangle$ from redshift $z=5$ onward. Given this, extrapolating the best fit of the $\langle f_{\rm esc}^{4\AA}\rangle$ (colored solid lines) to higher redshifts than $z=5$ is probably uncertain.
Moreover, we have estimated the ISM+CGM ${\rm Ly}\alpha$ escape fraction redshift evolution from the best fit of Hayes et al. (2011) global escape fraction and our IGM escape fraction. The mean ISM escape fraction, $f_{\rm esc}^{\rm ISM}$, is computed as $\langle f_{\rm esc}^{\rm Lya}\rangle/$$\langle f_{\rm esc}^{4\AA}\rangle$ and is displayed as the blue and green dotted line when using $\langle f_{\rm esc}^{4\AA}\rangle$ from IGM+z and IGM-z, respectively. The global ${\rm Ly}\alpha$ escape fraction is dominated by the ISM up to redshift $z=5$ when the ISM mean escape fraction matches $\langle f_{\rm esc}^{4\AA}\rangle$. Then, $f_{\rm esc}^{\rm ISM}$ grows until it reaches unity at $z=6$ (extrapolating the IGM+z and IGM-z best fits). If we consider that $\langle f_{\rm esc}^{4\AA}\rangle$ keeps constant from redshift $z=5$ onward, then the $f_{\rm esc}^{\rm ISM}$ would reach unity at $z=7$.
At the same time, our best fit for $\langle f_{\rm esc}^{4\AA}\rangle$ and for $f_{\rm esc}^{\rm Lya}$ by Hayes et al. (2011) cross at redshift 6 (or $z=7$ if the $\langle f_{\rm esc}^{4\AA}\rangle$ is kept constant from $z=5$). This is puzzling as the global ${\rm Ly}\alpha$ escape fraction should always be lower than that of the IGM component. Considering the case at high redshift in which the $f_{\rm esc}^{\rm ISM}=1$ due to the low dust abundance, only the IGM would obscure the ${\rm Ly}\alpha$ line. In this scenario, the global escape fraction should match that of the IGM.
We find some tension when extrapolating our results on the IGM escape fraction and those on the global escape fraction in the literature. On the one hand, Goovaerts et al. (2024)’s measurement at redshift $z=6$ suggests that the global ${\rm Ly}\alpha$ escape fraction could be compatible with unity already. At the same time, extrapolating Hayes et al. (2011), $f_{\rm esc}^{\rm Lya}$ would reach unity at $z\sim 11$. On the other hand, we find that the IGM escape fraction decreases with redshift up to $z=5.0$. Additionally, somewhere between redshift 5 and 11, the epoch of reionization takes place, meaning a higher abundance of neutral hydrogen in the universe, particularly in the IGM. This will cause a much lower $\langle f_{\rm esc}^{4\AA}\rangle$ if the typical ${\rm Ly}\alpha$ line profile remains the same, from $z=0$ to $z=6$.
This tension could be alleviated if the typical ${\rm Ly}\alpha$ line profile at $z>6$ is much more dominated by a red peak than those observed at lower redshifts. If the ${\rm Ly}\alpha$ radiation is redshifted enough within galaxies, then the IGM would have less impact. Thus, $\langle f_{\rm esc}^{4\AA}\rangle$ could increase at higher redshifts.
Another scenario is that neither extrapolating our estimation for the IGM escape fraction nor those like that of Hayes et al. (2011) is a valid option. Until now, the global ${\rm Ly}\alpha$ escape fraction has only been measured up to $z=6$. Measurements at higher redshifts can potentially further insights.
4.8 Stacked Lyman-alpha line profile fitting
@@
In this section, we analyze the redshift dependence of the observed ${\rm Ly}\alpha$ stacked line profile shape (a similar analysis was done by Hayes et al., 2021a, 2023; we focus here on the reconstruction of the pre-IGM spectra using zELDA). We limit our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$9.0. This ensures that more than $\sim$85of the ${\rm Ly}\alpha$ line profiles of HST are recovered with KS¡0.1. Meanwhile, as $W_{\rm in}$ changes from 1.25Å to 4Å in MUSE data, $\sim$85and $\sim$65of the line profiles should be recovered with KS¡0.1, respectively (see appendix D in ZP24). $S/N_{p}$9.0 was chosen as a compromise between reconstruction quality and number counts. We verify that similar results are obtained across different signal-to-noise thresholds ($S/N_{p}$6.0, 7.0, 8.0, 10.0), with higher thresholds producing noisier stacked line profiles, while the overall trends remain consistent.
@@
In Fig. 12, we present the ${\rm Ly}\alpha$ stacked line profiles across six redshift bins: $0.50.0<z<$ 0.5 (94 sources), $3.53.0<z<$ 3.5 (19 sources), and $4.03.5<z<$ 4.0 (17 sources) in the top row (left to right), and $4.54.0<z<$ 4.5 (10 sources), $5.04.5<z<$ 5.0 (13 sources), and $6.05.0<z<$ 6.0 (5 sources) in the bottom row (left to right). To ensure a fair comparison, the ${\rm Ly}\alpha$ profiles predicted by the ANN models are computed at the same resolution ($W_{\rm g}$) as the observations.
In general, we find that the observed stack spectrum (shown in grey in Fig. 12) changes with redshift. At $0.50.0<z<$ 0.5 we find a faint but detected continuum. Meanwhile, at higher redshift, no continuum is detected. At $0.50.0<z<$ 0.5, the ${\rm Ly}\alpha$ stacked line profile exhibits a clear double peak line profile, with a prominent red peak that extends $\sim 2\AA{}$ redder than ${\rm Ly}\alpha$. The blue peak is weaker and also extends $\sim 2\AA{}$ bluewards ${\rm Ly}\alpha$. The same structure of the double peak is also apparent at $3.53.0<z<$ 3.5. Thus, the ${\rm Ly}\alpha$ stacked line profiles for $0.50.0<z<$ 0.5 and $3.53.0<z<$ 3.5 exhibit only minor differences. At $4.03.5<z<$ 4.0, the blue peak becomes less pronounced and eventually disappears for $z>4.0$, while the red peak remains consistently present across all redshift bins. This is consistent with the findings by Hayes et al. (2021b).
Again, the observed stacked ${\rm Ly}\alpha$ line profile at redshift $0.50.0<z<$ 0.5 and that at $3.53.0<z<$ 3.5 exhibit a very similar blue peak. This is consistent with our measurements of the $\langle f_{\rm esc}^{4\AA}\rangle$ and the redshift invariance of the galactic stack line profile. We find that $\langle f_{\rm esc}^{4\AA}\rangle$ changes little from $z\sim 0.5$ to $z\sim 3.0$. Therefore, if the galactic stacked line profile is the same at $z\sim 0.5$ to $z\sim 3.0$, then the observed stack should be very similar at these redshifts. Reassuringly, this is what we find.
@@
Figure 13: Six examples of zELDA’s prediction on observed line profiles displayed in rest frame. The observed line profile is displayed on dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s galactic line profiles of IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source and the outflow properties for each models are shown in the corresponding color. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å.The central grey shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The red arrows show mark potential IGM features.
@@
In comparison, IGM+z and IGM-z behave similarly. Overall, the ${\rm Ly}\alpha$ stacked line profile predicted by both ANN models remains unchanged through redshift. Both ANN models predict that the red peak becomes slightly narrower with increasing $z$, as found in the observations. However, both IGM+z and IGM-z predict that the amplitude of the blue peak is relatively constant and that it becomes slightly narrower at higher redshift.

### v2

@@ 5.2 Line profile fitting accuracy
Figure 8: Comparison between stacked ${\rm Ly}\alpha$ line profiles. The observed stacked ${\rm Ly}\alpha$ line profile is shown in grey, while zELDA’s prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The stacked ${\rm Ly}\alpha$ line profile is computed in 6 redshift bins. The relevant redshift interval.
[…]
In Fig. 7, we show the distribution of the KS estimator for the IGM+z (blue), IGM-z (green), and NoIGM (yellow) in the red tails of the HST sources. We find that IGM-z and IGM+z outperform NoIGM. In particular, the median of the KS distribution for the models is 0.07 for IGM+z, 0.07 for IGM-z, and 0.09 for NoIGM. Meanwhile, NoIGM exhibits a little bit more dispersion than IGM+z and IGM-z. We conclude that the IGM+z and IGM-z models predict the red tail of the observed ${\rm Ly}\alpha$ line profile with the same precision.
5.3 Stacked Lyman-alpha line profile
@@
In this section, we analyze the redshift dependence of the observed ${\rm Ly}\alpha$ stacked line profile shape (a similar analysis was done by Hayes et al., 2021a, 2023; we focus here on the reconstruction of the pre-IGM spectra using zELDA). We limit our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$ >>9.0. This ensures that more than $\sim$85% of the ${\rm Ly}\alpha$ line profiles of HST are recovered with KS<<0.1. Meanwhile, as $W_{\rm in}$ changes from 1.25Å to 4Å in MUSE data, $\sim$85% and $\sim$65% of the line profiles should be recovered with KS<<0.1, respectively (see appendix D in Z25). $S/N_{p}$ >>9.0 was chosen as a compromise between reconstruction quality and number counts. We verify that similar results are obtained across different signal-to-noise thresholds ($S/N_{p}$ >>6.0, 7.0, 8.0, 10.0), with higher thresholds producing noisier stacked line profiles, while the overall trends remain consistent.
@@
In Fig. 8, we present the ${\rm Ly}\alpha$ stacked line profiles across six redshift bins: $0.0<z<0.5$ (94 sources), $3.0<z<3.5$ (19 sources), and $3.5<z<4.0$ (17 sources) in the top row (left to right), and $4.0<z<4.5$ (10 sources), $4.5<z<5.0$ (13 sources), and $5.0<z<6.0$ (5 sources) in the bottom row (left to right). To ensure a fair comparison, the ${\rm Ly}\alpha$ profiles predicted by the ANN models are computed at the same resolution ($W_{\rm g}$) as the observations.
In general, we find that the observed stack spectrum (shown in grey in Fig. 8) changes with redshift. At $0.0<z<0.5$ we find a faint but detectable continuum. Meanwhile, at higher redshifts, no continuum is detected. At $0.0<z<0.5$, the ${\rm Ly}\alpha$ stacked line profile exhibits a clear double peak line profile, with a prominent red peak that extends $\sim 2\AA$ {} redder than ${\rm Ly}\alpha$. The blue peak is weaker and also extends $\sim 2\AA$ {} bluewards ${\rm Ly}\alpha$. The same structure of the double peak is also apparent at $3.0<z<3.5$. Thus, the ${\rm Ly}\alpha$ stacked line profiles for $0.0<z<0.5$ and $3.0<z<3.5$ exhibit only minor differences. At $3.5<z<4.0$, the blue peak becomes less pronounced and eventually disappears for $z>4.0$, while the red peak remains consistently present across all redshift bins. This is consistent with the findings by Hayes et al. (2021b), i.e., these results are compatible with the scenario in which the ${\rm Ly}\alpha$ line profile is attenuated by the IGM progressively at higher redshifts and the pre-IGM spectra show little evolution. The IGM absorption at these redshifts mainly affects the blue side of ${\rm Ly}\alpha$ emission. This causes the main evolution in the ${\rm Ly}\alpha$ stacked line profile to be at the blue peak.
Again, the observed stacked ${\rm Ly}\alpha$ line profile at redshift $0.0<z<0.5$ and that at $3.0<z<3.5$ exhibit a very similar blue peak. This is consistent with our measurements of the $\langle f_{\rm esc}^{4\AA }\rangle$ and the redshift invariance of the galactic stack line profile. We find that $\langle f_{\rm esc}^{4\AA }\rangle$ changes little from $z\sim 0.5$ to $z\sim 3.0$. Therefore, if the galactic stacked line profile is the same at $z\sim 0.5$ to $z\sim 3.0$, then the observed stack should be very similar at these redshifts. Reassuringly, this is what we find. This would imply that the IGM absorption in the ${\rm Ly}\alpha$ line profiles would not evolve strongly from $z\sim 0$ to $z\sim 3$. In contrast, the mean IGM transmission changes significantly in this redshift range. $\langle T_{\rm IGM}\rangle$ evolves from $\sim 1$ at $z=0$ to $\sim 0.75$ at $z=3$ (Faucher-Giguère et al., 2008). Thus, our results would indicate that the stacked ${\rm Ly}\alpha$ line profile is not very sensitive to the IGM evolution in this redshift window.
@@
In comparison, IGM+z and IGM-z behave similarly. Overall, the ${\rm Ly}\alpha$ stacked line profile predicted by both ANN models remains unchanged through redshift. Both ANN models predict that the red peak becomes slightly narrower with increasing $z$, as found in the observations. However, both IGM+z and IGM-z predict that the amplitude of the blue peak is relatively constant and that it becomes slightly narrower at higher redshift. This suggests that the physical mechanisms shaping the ${\rm Ly}\alpha$ line profile do not evolve strongly through cosmic time.
In summary, our reconstruction of the pre-IGM spectra using stacks strongly supports the picture suggested by Hayes et al. (2021b) in which the evolution of the observed ${\rm Ly}\alpha$ line profile shape is mainly due to the IGM. This further validates our pre-IGM line profile reconstruction. While they use a similar dataset (obtained from LASD; Runnholm et al., 2021), the method is very different from ours. While they construct a mean IGM transmission from empirical models, we made our analysis on a basis of source per source procedure. The strength of our methodology in comparison to theirs is that zELDA provides the reconstructed spectrum of individual sources affected by very different IGM line of sights.
5.4 Comparison between $\langle f_{\rm esc}^{4\AA }\rangle$ and the IGM mean transmission
Figure 9: Left: illustration of the Lya emission line IGM escape fraction computation from the stacked line profile. The solid colored lines show the mean IGM transmission of Byrohl & Gronke (2020) in increasing redshift from purple to red. These curves are recalibrated to match the mean IGM transmission of Faucher-Giguère et al. (2008). The dashed green and blue line show the reconstructed stack line profiles using zELDA models (IGM+z in blue and IGM-z in green). The shaded region marks the wavelength window there the Lya emission line IGM escape fraction is computed. Right: Mean Lya IGM escape fraction as a function of redshift (squares) for IGM+z (blue). The solid lines indicate the fits, including all the sources. Meanwhile, the colored dashed line indicates the fit using only MUSE data. The grey dashed line shows the Faucher-Giguère et al. (2008) mean IGM transmission. The yellow circles show the mean Lya IGM escape fraction using the Faucher-Giguère et al. (2008) recalibrated IGM transmission curves applied to the reconstructed stacked spectrum (i.e., convolving the solid lines with the dashed lines from the left panel).
The connection between the average IGM transmission, $\langle T_{\rm IGM}\rangle$, and the mean IGM ${\rm Ly}\alpha$ escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, is initially unknown. $f_{\rm esc}^{4\AA }$ quantifies the fraction of ${\rm Ly}\alpha$ radiation escaping from LAEs over intermediate IGM scales (few cMpc), after being perturbed by RT in the ISM and CGM. $\langle T_{\rm IGM}\rangle$ is derived from the spectra of extragalactic sources within a wavelength window located several angstroms blueward of the ${\rm Ly}\alpha$ line, measuring the IGM hundreds of cMpc from the source. As a first approximation, the IGM primarily affects the blue side of the ${\rm Ly}\alpha$ line. Consequently, the measurement of $f_{\rm esc}^{4\AA }$ around the ${\rm Ly}\alpha$ wavelength naturally distinguishes $\langle T_{\rm IGM}\rangle$ from $\langle f_{\rm esc}^{4\AA }\rangle$. In addition, $f_{\rm esc}^{4\AA }$ is obtained by convolving the ${\rm Ly}\alpha$ line profile with the IGM transmission along a sightline. Therefore, the shape of the ${\rm Ly}\alpha$ line profile is crucial for assessing $\langle f_{\rm esc}^{4\AA }\rangle$, whereas $\langle T_{\rm IGM}\rangle$ remains independent of it. For instance, given a constant IGM transmission, a line profile with a stronger red peak than blue peak yields a higher $f_{\rm esc}^{4\AA }$, and vice versa.
To compare our measurement of $\langle f_{\rm esc}^{4\AA }\rangle$ with previous works, we compute a synthetic estimate of $\langle f_{\rm esc}^{4\AA }\rangle$ based on the redshift evolution of $\langle T_{\rm IGM}\rangle$ from Faucher-Giguère et al. (2008). Specifically, we calculate the mean ${\rm Ly}\alpha$ escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, that corresponds to the mean IGM absorption reported by Faucher-Giguère et al. (2008), by convolving the mean IGM transmission curves around ${\rm Ly}\alpha$ with the average ${\rm Ly}\alpha$ line profile as a function of redshift. For this, we adopt the IGM transmission shape from Byrohl & Gronke (2020), recalibrated to match Faucher-Giguère et al. (2008). For the mean ${\rm Ly}\alpha$ line profile emerging from the ISM/CGM, we use the reconstructed stacked profile discussed in 5.3, which includes all sources from $z=0$ up to $z=5$. We repeated the analysis using stacked profiles in redshift bins and obtained consistent results, as the stacked spectrum varies little with redshift.
In the left panel of Fig. 9, we show the mean IGM transmission curves as a function of redshift (solid colored lines). The reconstructed stacked spectrum between $z=0$ and $z=5$, as modeled by IGM+z and IGM-z, is displayed in dashed blue and green lines, respectively. We compute the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ as the ratio between the convolution of the mean IGM transmission and the stacked line profile, and the unabsorbed stacked profile, within a window of $\pm 2\AA$ {} around ${\rm Ly}\alpha$ (grey region).
The synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ derived from the stacked line profile and mean IGM transmission curves is shown as yellow dots in the right panel of Fig. 9. The synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ decreases with redshift because the stacked line profile emerging from the ISM/CGM remains constant, while the blue peak is increasingly absorbed at higher $z$ as $\langle T_{\rm IGM}\rangle$ decreases (dashed grey curve). We find excellent agreement between the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ and our measurements from the ${\rm Ly}\alpha$ line profiles (blue squares), for both the IGM+z and IGM-z (not shown) models. At $z>2$, the synthetic and measured $\langle f_{\rm esc}^{4\AA }\rangle$ values are consistent within $1\sigma$, indicating good agreement with $\langle T_{\rm IGM}\rangle$. At $z<1$, however, we find that the measured $\langle f_{\rm esc}^{4\AA }\rangle$ is $\sim 8\%$ lower than the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ obtained using Faucher-Giguère et al. (2008).
Considering the parametric estimation of the redshift evolution of $\langle f_{\rm esc}^{4\AA }\rangle$, we find that when only MUSE data are used (dashed blue line), there is very good agreement with the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$, showing consistency with the $\langle T_{\rm IGM}\rangle$ evolution from Faucher-Giguère et al. (2008). However, when including the HST dataset, the parametric estimate lies $\sim 8\%$ below the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ at $z<1.0$, while remaining consistent at $z>2$.
This result indicates that zELDA predicts ${\rm Ly}\alpha$ line profiles to be more absorbed than expected given the mean IGM optical depth at $z<1$, whereas at $z>1$ the predictions match the literature $\langle T_{\rm IGM}\rangle$ values. In Z25, we tested biases and accuracy in recovering $\langle f_{\rm esc}^{4\AA }\rangle$ as a function of redshift, showing that both IGM+z and IGM-z correctly recover $\langle f_{\rm esc}^{4\AA }\rangle$ at low redshift for a range of $\langle f_{\rm esc}^{4\AA }\rangle$ $(z)$ relations. Therefore, our findings may indicate either a limitation of the shell model at low $z$, or that ${\rm Ly}\alpha$ line profiles are indeed more attenuated than expected at $z<0.5$. In Sect. 5.6, we show several examples of $z<0.5$ sources that may exhibit IGM-attenuated ${\rm Ly}\alpha$ spectra, supporting this interpretation.
For comparison, we also show $\langle T_{\rm IGM}\rangle$, the mean IGM transmission from (Faucher-Giguère et al., 2008), as a dashed grey line. At $z<0.5$, the IGM-z and IGM+z predictions of $\langle f_{\rm esc}^{4\AA }\rangle$ fall $\sim 0.1$ below $\langle T_{\rm IGM}\rangle$. The difference between $f_{\rm esc}^{4\AA }$ and $\langle T_{\rm IGM}\rangle$ can be attributed to stronger absorption around the ${\rm Ly}\alpha$ line center (Laursen et al., 2011b; Gurung-López et al., 2022). Even if the IGM transmission is close to unity a few angstroms away from ${\rm Ly}\alpha$, stronger absorption in the range $2\AA {}<\Delta\lambda<0.5\AA {}$ can affect the blue peak of $z=0$ lines. At higher redshift ($z>2.0$), both the best-fit parametric curves for the ANN models and the predicted $\langle f_{\rm esc}^{4\AA }\rangle$ lie above $\langle T_{\rm IGM}\rangle$. This is because most of the ${\rm Ly}\alpha$ flux resides in the red peak of the ${\rm Ly}\alpha$ line profile, which remains largely unaffected by the IGM up to $z=5.0$. Thus, although $\langle T_{\rm IGM}\rangle$ can be as low as $\sim 10\%$ at $z=5.0$, a large fraction of ${\rm Ly}\alpha$ photons still escape through the IGM, as they are already redshifted relative to the ${\rm Ly}\alpha$ wavelength by the ISM/CGM. However, this trend could also reflect a selection effect: in this work, we measure $f_{\rm esc}^{4\AA }$ in MUSE galaxies with strong observed ${\rm Ly}\alpha$ emission rather than in the full galaxy population. Consequently, flux-limited samples may yield $\langle f_{\rm esc}^{4\AA }\rangle$ values higher than those of the overall star-forming galaxy population.
5.5 Comparison between the ${\rm Ly}\alpha$ global and IGM escape fractions
Figure 10: Comparison between the redshift evolution of the IGM-only and global ${\rm Ly}\alpha$ escape fractions. The filled colored squares show the IGM ${\rm Ly}\alpha$ escape fraction ($\langle f_{\rm esc}^{4\AA }\rangle$) found by IGM+z up to $z=5.0$, while those empty show the results at $z>5$. The colored solid line show the best fit to $\langle f_{\rm esc}^{4\AA }\rangle$ as presented previously. The dashed red line shows the global ${\rm Ly}\alpha$ escape fraction estimation by Hayes et al. (2011). The grey symbols show different studies from the literature measuring the global ${\rm Ly}\alpha$ escape fraction and their uncertainty.
Figure 11: Six examples of zELDA’s prediction on observed line profiles displayed in rest frame. The observed line profile is displayed in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s galactic line profiles of IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source and the outflow properties for each model are shown in the corresponding color. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å.The central grey shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The red arrows mark potential IGM features.
[…]
In Fig. 10, we show the $f_{\rm esc}^{\rm Lya}$ redshift evolution obtained by Hayes et al. (2011) (dashed red line) along with the individual measurements from the other studies in grey as described in the legend. Other studies, such as Goovaerts et al. (2024) (based on Thai et al. (2023)), also obtained relations between the redshift and $f_{\rm esc}^{\rm Lya}$ , finding compatible results with Hayes et al. (2011). Overall, it is found in the literature that $f_{\rm esc}^{\rm Lya}$ increases with redshift, being earlier galaxies less attenuated in ${\rm Ly}\alpha$. Normally, this is attributed to the dust in the ISM of galaxies, which is less abundant at high redshift (Hayes et al., 2011; Gurung-López et al., 2019b; Goovaerts et al., 2024). In particular, the different experiments find that $f_{\rm esc}^{\rm Lya}<5\%$ at $z<3.5$. Also, at $5<z<6$, Ouchi et al. (2008); Cassata et al. (2011); de La Vieuville et al. (2019) found a $f_{\rm esc}^{\rm Lya}$ below 0.2 with relatively low uncertainty of $\sim 0.1$. Meanwhile, with larger uncertainties, Sobral et al. (2018) found a $f_{\rm esc}^{\rm Lya}\sim 0.3$ at $4<z<6$ and Goovaerts et al. (2024) found $f_{\rm esc}^{\rm Lya}\sim 0.8$ at $z=6$. Given the uncertainties, all these studies are mutually compatible.
[…]
Moreover, we find that the IGM contributes only marginally to the global ${\rm Ly}\alpha$ escape fraction up to $z\sim 3$, where $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.8$ while the total $f_{\rm esc}^{\rm Ly\alpha}$ remains below 0.1. At higher redshift, as the IGM becomes increasingly dense, its role in regulating the global escape fraction becomes more significant. In particular, at $z=5.0$ zELDA predicts $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.5$, while $f_{\rm esc}^{\rm Ly\alpha}\sim 0.2$. This implies that at this epoch the ISM/CGM absorbs about 40% of the ${\rm Ly}\alpha$ radiation. Thus, our results indicate that at redshift 5.0 the IGM is already the main absorber of the ${\rm Ly}\alpha$ emission line.
Our results therefore suggest that the IGM escape fraction decreases steadily with redshift, at least up to $z=5$. If this trend continues beyond $z=5$, the IGM would become more dominant, controlling the global ${\rm Ly}\alpha$ escape fraction, while the relative impact of the ISM/CGM would diminish.